Problem 6-47 Present Value and Multiple Cash Flows [LO1] What is the value today of $4,700 per year, at a discount rate of 9 percent, if the first payment is received 5 years from today and the last payment is received 15 years from today? (Do not round intermediate calculations and round your final answer to 2 decimal places, e.g., 32.16.)
Answer #1

The 4700 is the annuity due payment at the beginning period of 6th year.

Present value of future cash inflows = amount * Present value factor at 9% for ten years

= 4700*6.9952 = 32877.44

32877.44 is the value of money received at the beginning of the 6th year.

The present value of 32877.44 on the year 1 = Future value * present value factor for a lump sum

= 32877.44 *0.6499 = 21367.048

The Present value today in year one 21367.048